# TuneShield: Mitigating Toxicity in Conversational AI while Fine-tuning on Untrusted Data

### v1

@@ VI-F What level of bias in the toxicity classifier can TuneShield tolerate?
Table VI: RTR for LLaMA-2 chatbot fine-tuned with TuneShield, with varying Recall for toxic sub-categories within specialized category while maintaining default Recall for other sub-categories.
VI-G Comparison with existing alignment schemes
We compare TuneShield with existing safety alignment schemes that can mitigate toxicity in chatbots.
@@ VI-G Comparison with existing alignment schemes
Use of LLMs as training data filters to mitigate toxicity.    He et al. [29] uses prefix tuning, a supervised learning scheme to adapt an LLM for toxicity detection. Prefix tuning involves updating trainable vectors called prefixes while keeping the LLM’s parameters frozen. Although He et al. [29] shows better transferability across multiple datasets, their best transferable model (T5 base fine-tuned on MHS [37]) achieves F1 scores of 48.86% and 60.23% on our offensive and specialized datasets, respectively, which are significantly lower than TuneShield’s LLM-based classifiers.
HateGuard [73] uses Chain-of-Thought (CoT) prompting to detect identity hate speech, i.e., hate speech directed or incited against individuals or groups in online social media. This is a narrower focus than our broader, identity-agnostic toxicity definition. We carefully analyzed HateGuard to compare it with our work but were unsuccessful due to several problems we discovered, including incorrect data labels, test data snooping, and other issues. Details are in the Appendix -G. We have communicated our findings with the authors of HateGuard.
Concurrent work by Hu et al. [32] also uses LLM refusal tokens for toxicity classification, similar to our refusal. However, their detector requires labeled data to establish a decision boundary, unlike our unsupervised approach.
Alignment via baking in awareness of toxic language during training.    Rather than filtering toxic samples from the training set, these defenses aim to instill knowledge of toxic language during training, to steer the chatbot away from toxic responses [24, 8, 4]. These methods use toxicity classifiers to "bake in" awareness by labeling samples in the training data as toxic or non-toxic. Gehman et al. [24] and Baheti et al. [8] use Attribute Conditioning (ATCON) [38], training models with control tokens (<toxic>, <non-toxic>) and generating responses conditioned on a non-toxic token. We do not compare with ATCON, as Weeks et al. [78] found it ineffective against toxicity injection. Cringe Loss [4] penalizes toxic token generation during training but requires a labeled dataset for detoxification. ADPO [39], which refines ATCON using direct preference optimization (DPO) [60], relies on a labeled dataset of toxic and non-toxic utterances from the victim chatbot itself and the ability to controllably generate it. In contrast, our method has no such data requirement.
Generation steering defenses.   These are inference-time defenses applied to steer the trained chatbot away from generating toxic responses. These approaches rely on the feedback from a toxicity classifier or other LMs to influence response generation [24, 48, 41, 58, 51, 66]. While these defenses can supplement our framework, they do not address toxicity learned during fine-tuning. This remains a challenge when the chatbot is further fine-tuned downstream.
VII TuneShield: Adversarial Robustness
@@
We study the adversarial robustness of TuneShield considering an adaptive attacker who aims to poison the fine-tuning dataset with “adversarial” toxic samples. The adversarial samples are crafted to bypass the guard rails provided by TuneShield. We study attacks targetting two core components of TuneShield—toxicity classifier and healing data generator. Additionally, we assess TuneShield’s ability to mitigate adversarial attacks within a dialog-based learning (DBL) setting.
VII-A Attacks against toxicity classification
@@
Adversaries craft toxic samples to be misclassified as non-toxic, ensuring these samples remain as part of the training dataset. If a significant portion of the adversarial samples can fool the toxicity classifier, the rest of TuneShield’s pipeline will fail to mitigate toxicity. Recall that the second stage of TuneShield relies on toxicity classifiers to create healing and preference data for fine-tuning and alignment. We study two types of attacks against toxicity classifiers—a traditional adversarial attack, and jailbreak attacks.
VII-A.1. Adversarial attack
@@
Most prior work on fooling text classifiers targets traditional NLP systems [36], and not LLM-based classifiers. In fact, recent work argues that state-of-the-art adversarial attacks targeting traditional NLP classifiers are not effective against LLM-based models [13]. Therefore, we use PromptAttack [84], a recent approach designed specifically to fool LLM-based classifiers, shown to outperform other attacks [76, 75, 45, 85]. We consider a gray-box setting, where the adversary is unaware of the exact LLM used by the defender but has knowledge of the prompts employed for toxicity classification. The adversary has no query access to the victim classifier.
@@
Given a context-response sample $x$, PromptAttack uses an off-the-shelf LLM, $LLM_{p}$, to perturb the sample to make it adversarial, i.e., $x_{adv}$. This is done using a novel prompt-based strategy. This prompt has three parts: (a) Input sample, $x$. (b) Attack objective: an instruction to fool a surrogate LLM-based toxicity classifier, $LLM_{s}$, while maintaining semantic similarity between $x$ and $x_{adv}$. PromptAttack recommends using the same model for $LLM_{p}$ and $LLM_{s}$. (c) Attack guidance: instructions to perturb the sample $x$ using 9 different strategies at the character, word, and sentence level. Finally, PromptAttack uses an ensemble strategy that generates multiple adversarial samples using the different perturbation strategies. The adversarial sample that successfully fools the surrogate $LLM_{s}$ and has the highest semantic similarity is eventually chosen. We use the LLaMA-2-Chat 13B as $LLM_{p}$ and 7B as $LLM_{s}$ model [^5] LLaMA-2-Chat 7B gave better attack performance than 13B model.. Details are in Appendix -F.
@@
TuneShield (NH)  0.2 6.17 0.109   1.8 6.23 0.103
Table VII: RTR and utility metrics for LLaMA-2 chatbot fine-tuned with TuneShield for specialized category under adversarial attack targeting refusal classifier. $\Delta$R shows Recall degradation for the toxic class. Adv. Attack row presents adversarial attack targeting the refusal (Ref.) classifier without TuneShield.
Table VII shows results for TuneShield using the refusal classifier (NH setting) on the challenging LLaMA-2 chatbot. For the offensive and specialized categories, our adversarial samples successfully inject toxicity (RTR $>47\%$), while preserving semantic similarity. Our TuneShield defense effectively mitigates toxicity in both categories (RTR $<1.8\%$), significantly lower than no-attack. Upon further investigation, we find that PromptAttack did not produce highly effective adversarial samples. For example, in offensive, there was only a 2.86% degradation in Recall ($\Delta R$). Table XII (Appendix) has results for other configurations of TuneShield, and we observe similar successful outcomes.
VII-A.2. Jailbreak attacks
Jailbreak attacks bypass LLM safety mechanisms to generate otherwise restricted responses [87]. We leverage existing work on jailbreak attacks to fool our refusal classifier. In our setting, the attacker can only add the jailbreak instructions/prompts to the context-response pairs (i.e., the poisoned samples), and cannot modify the prompt used by our refusal classifier. We consider a more challenging setting for the defender—the attacker has full knowledge and access to our refusal classifier, a white-box attack. We investigate 2 jailbreak attacks: (1) Manually-designed jailbreak attack and (2) Optimization-based attack.
@@ VII-A.2. Jailbreak attacks
Table VIII shows the results for LLaMA-2 chatbot using TuneShield (using CH setting for specialized category). The manually-designed jailbreak attack was ineffective against TuneShield. Instead of degrading classifier performance, it increased Recall by 3.21%. This is due to sandwich prevention in our refusal classifier’s prompt, where the instruction is placed at the end after the data sample—a technique shown to be an effective defense in prior work [50]. [^6] Without sandwich prevention, the Recall would degrade by 37.88%. Subsequently, applying TuneShield on this attack significantly reduces RTR from 58.3% (attack setting) to 4.5% (below no-attack level).
Optimization-based jailbreak attack.    This attack optimizes an adversarial suffix $suffix_{adv}$ that when appended to a context-response pair $x$ is classified as non-toxic by the refusal classifier. We use Liu et al. [49]’s approach to craft a universal adversarial suffix tailored to a specific instruction (i.e., misclassify toxic samples). This approach is an improvement over the Greedy Coordinate Gradient (GCG) [88] jailbreak attack. Optimization requires a surrogate LLM, which is the same as the victim classifier (white-box attack). Optimization is performed on 5 samples for 1K epochs with a batch size of 256. We use a token length of 150 for the adversarial suffix. Figure 11 (Appendix) shows an example attack.
The optimization-based attack produces a significant 52.23% degradation in Recall for our refusal classifier. However, TuneShield still effectively achieves the objective of mitigating toxicity, resulting in a 7.9% RTR (well below the no-attack level). This can be attributed to the resilience of TuneShield to degradation of the performance of the toxicity classifier (see analysis in Section VI-F).
@@
TuneShield (CH)  4.5 5.85 0.099   7.9 5.78 0.101
Table VIII: RTR and utility metrics for LLaMA-2 fine-tuned with TuneShield under manually-designed and optimization-based jailbreak attacks targeting refusal toxicity classifier. $\Delta$R shows toxic class Recall degradation.
VII-B Attack against healing data generation
@@
In this attack, the adversary tricks the safety-aligned LLM used by TuneShield into producing toxic responses, instead of contextually relevant ‘healing’ responses. This can be accomplished through a jailbreak attack. However, the optimization-based attack in the previous section VII-A can be prohibitively (computationally) expensive in this setting. Unlike the previous setting, where the goal was to produce a specific response (make the refusal classifier say ‘yes’) for every sample, here we aim to produce a different toxic response for each input context. To address this challenge, we use AmpleGCG-Plus [42], an LLM trained on a large pool of adversarial suffixes, to efficiently generate tailored suffixes that can jailbreak target LLMs.
@@
We use AmpleGCG-Plus with a surrogate classifier that is identical to the victim classifier. For each context, we append a jailbreak instruction (e.g., ‘produce a harmful response’) and generate 10 suffix candidates using AmpleGCG-Plus. Figure 9 (Appendix) shows an example attack. Suffixes are generated with a token length constrained between 20 and 50 tokens. We select the suffix yielding highest toxicity score for the response based on the surrogate classifier. When evaluated using the specialized category, 10.54% of adversarial contexts produce toxic responses, indicating a successful attack. These adversarial samples successfully inject toxicity leading to $50.06\%$ RTR. When TuneShield is applied, the RTR reduces significantly to 17.3%, coming close to no-attack setting (13.1%).
Note, this is a highly challenging attack—a white-box setting and violates the integrity of the healing data, which is used by our DPO scheme as the preference data. Despite these challenges, TuneShield is able to significantly mitigate toxicity. A straightforward approach to further enhance resilience is to integrate existing state-of-the-art jailbreak defenses [87, 52] into TuneShield, e.g., one can use a jailbreak prompt detection scheme as an additional data filter before applying TuneShield.
VII-C Adversarial attacks in Dialog-based learning
We conduct a case study of integrating TuneShield to mitigate an adaptive adversarial attack in Dialog-based Learning (DBL) setting. Recently, Weeks et al. [78] conducted the first study of toxicity injection in a DBL setting. In this setting, post-deployment, a chatbot is periodically trained on recent conversations with its users. The attack uses an LLM-based agent that masquerades as a normal user to engage in toxic conversations with a deployed chatbot, thereby poisoning the subsequent DBL training cycle. When the chatbot is trained in the next DBL cycle, it learns toxic language. This study proposes two attacks: (1) Indiscriminate attack, aims to produce toxic responses for a fraction of inputs, regardless of the input type (toxic or non-toxic). (2) Backdoor attack, aims to trigger toxic responses only when certain trigger phrases are present in the input. We consider their most challenging adversarial attack setting, i.e., an adaptive attacker who aims to bypass toxicity filters. This is an excellent setting to evaluate TuneShield for multiple reasons: (1) its practical relevance (toxicity injection post-deployment), (2) studies multiple injection strategies, and (3) uses a toxic language distribution different from our previous evaluation.
We obtained code and datasets from Weeks et al. to reproduce their attacks. The base model is a DD-BART model, i.e., a BART model fine-tuned on the DailyDialog [46] dataset. To create the victim chatbot, the base model is fine-tuned on simulated conversations between LLM-based agents and the deployed chatbot. Table IX shows the results when applying TuneShield to secure a DBL training cycle (averaged over 5 trials). RTR is computed separately for non-toxic and toxic/trigger inputs, as in the original work. We use the refusal classifier for TuneShield. For comparison, we include the Multi-level toxicity classifier from Weeks et al. [82], which filters toxic samples during training and inference. The full TuneShield pipeline using CH, successfully mitigates toxicity resulting in RTR close to 0 in most cases. TuneShield also outperforms the existing Multi-level filter. This also shows that TuneShield can provide protection against multiple attack strategies, i.e., indiscriminate and backdoor toxicity injection attacks. More experimental details are in Appendix -E.

### v2

@@ 6.6. What level of bias in the toxicity classifier can Optimus tolerate?
Table 5. RTR for LLaMA-2 chatbot fine-tuned with Optimus, with varying Recall for toxic sub-categories in specialized category (maintaining default Recall for other sub-categories).
6.7. Comparison with other defenses
6.7.1.. Using advanced filters to mitigate toxicity.
An obvious approach to mitigate toxicity is to use an ML-based toxic language classifier to filter toxic conversations from the fine-tuning dataset. In our analysis (Section 6.3), we have already demonstrated the limitations of such an approach. We still conduct empirical analysis of a recent representative approach.
@@ 6.7. Comparison with other defenses
LLM-based filter by He et al. (He et al., 2023):   This approach uses prefix tuning, a supervised learning scheme to adapt an LLM for toxicity detection. Prefix tuning involves updating trainable vectors called prefixes while keeping the LLM’s parameters frozen. Although He et al. (He et al., 2023) shows better transferability across multiple datasets, their best transferable model (T5 base fine-tuned on MHS dataset) achieves F1 scores of 48.86% and 60.23% on our offensive and specialized datasets, respectively, which are significantly lower than Optimus’s LLM-based classifiers. Supervised toxicity classifiers generalize poorly to toxic language distributions that differ from those in the training dataset. These results further highlight the limitations of a filtering-only approach.
Other filtering approaches: HateGuard (Vishwamitra et al., 2024) is another LLM-based filter that uses Chain-of-Thought (CoT) prompting to detect hate speech on Twitter. This work targets a highly specific type of toxicity on platforms like Twitter, limiting its applicability in practice. Concurrent work by Hu et al. (Hu et al., 2024) also uses LLM refusal tokens for toxicity classification, similar to our refusal. However, their detector requires labeled data to establish a decision boundary, unlike our unsupervised approach. Most other filtering schemes also rely on supervised learning, assuming access to a labeled dataset with a distribution similar to the target toxic language (Sun et al., 2021) e.g., BERT-based toxicity classifier (Hanu, 2021). This is not a practical assumption. Prior work has studied toxicity classifiers in a non-conversational setting, i.e., determining whether a given statement is toxic (per, 2023; Schick et al., 2021).
6.7.2.. Baking-in safety during fine-tuning.
Instead of simply filtering out toxic samples from the training data, these defenses modify the fine-tuning process itself to incorporate safety signals or constraints. Optimus falls in this category. We compare Optimus with the state-of-the-art scheme in this category called StarDSS (Peng et al., 2025).
StarDSS (Peng et al., 2025):    StarDSS steers the model towards safe behavior during fine-tuning via a novel dynamic safety shaping (DSS) scheme. While most existing defenses employ static safety shaping, where an entire training sample is atomically filtered or weighted based on a single safety score—StarDSS addresses the “blind spots” where safety context shifts within a single response. Specifically, it repurposes off-the-shelf guardrail models to score partial responses, generating a token-level signal termed the Safety Trajectory Assessment of Response (STAR). These fine-grained scores track the evolution of risk segment-by-segment, allowing the training objective to be dynamically re-weighted: learning is reinforced for safe segments while simultaneously suppressed for unsafe tokens.
We evaluate StarDSS on the specialized dataset using the recommended hyperparameters (chunk size of 5 and KL loss scaling of 0.5). StarDSS yields a high RTR of 54.3%, failing to satisfy the defense requirement of an RTR $\leq$ 13.1%. In contrast Optimus demonstrates superior mitigation in this challenging setting (Section 6.5). We attribute StarDSS’s underperformance to its reliance on the underlying guardrail model; the LlaMA-3 guardrail struggled to accurately score the specific toxic segments in our specialized benchmark. Furthermore, we observed significant training instability. Results are reported after averaging over two trials; the other failed to converge due to training instability, and even the successful run exhibited severe utility degradation with a perplexity (PPL) of ~229. We hypothesize that this instability stems from optimization conflicts between the aggressive token-level safety re-weighting and the KL divergence penalty intended to regularize the model. This comparative analysis further highlights the fundamental limitations of solely relying on filters, as their errors can lead to poor toxicity mitigation.
Other fine-tuning stage defenses: We omit comparisons with Lisa (Huang et al., 2024b) and Deep Token (Qi et al., 2025) as they are outperformed by StarDSS. We also exclude Safely Partial-Parameter Fine-Tuning (SPPFT) (Li et al., 2025c), which proposes freezing the gradients of “safety layers” to preserve alignment. We were unable to reproduce key results for SPPFT due to missing data in the provided artifacts. While SaLoRA (Li et al., 2025b) and LARF (Li et al., 2025a) propose safety-aware adaptation and filtering, respectively, both rely on external reference safety datasets. Furthermore, SaLoRA is restricted to LoRA fine-tuning. In contrast, Optimus is broadly applicable and mitigates toxicity without requiring additional reference data.
6.7.3.. Other defenses.
A pre-fine-tuning defense modifies the model initialization to induce safety robustness, ensuring the model resists unlearning its safety constraints even when subsequently fine-tuned on harmful data. This includes Booster (Huang et al., 2025b), Vaccine (Huang et al., 2024a). We omit these because StarDSS (our baseline) reportedly outperforms Vaccine, and because they address a different threat model where the attacker controls the full training process. In contrast, we assume a trustworthy pipeline where only the dataset is compromised. Post-fine-tuning defenses intervene after adaptation through inference-time steering (Gehman et al., 2020; Liu et al., 2021) or model repair (Huang et al., 2025a). Steering methods remain bottlenecked by the accuracy of their toxicity classifiers. Repair schemes such as Antidote (Huang et al., 2025a), which prune parameters using reference safety data, are reactive and suffer from two key limitations: (1) significant utility degradation, and (2) brittleness to distribution shifts when the reference data fails to cover the specific toxicity acquired during fine-tuning.
7. Resilience against Adaptive Attacks
@@
We consider an adaptive attacker who aims to poison the fine-tuning dataset with “adversarial” toxic samples. The adversarial samples are crafted to bypass the guard rails provided by Optimus. We study attacks targeting two core components of Optimus—toxicity classifier and healing data generator. Additionally, we assess Optimus’s ability to mitigate adversarial attacks within dialog-based learning (DBL).
7.1. Attacks against toxicity classification
@@
Adversaries craft toxic samples to be misclassified as non-toxic, ensuring these samples remain as part of the training dataset. If a significant portion of the adversarial samples fool the toxicity classifier, Optimus’s pipeline will fail to mitigate toxicity. Recall that the second stage of Optimus relies on toxicity classifiers to create healing and preference data for fine-tuning and alignment. We study two types of attacks against toxicity classifiers—a traditional adversarial attack, and jailbreak attacks.
7.1.1.. Adversarial attack
@@
Most prior work on fooling text classifiers targets traditional NLP systems (Jin et al., 2020), and not LLM-based classifiers. In fact, recent work argues that state-of-the-art adversarial attacks targeting traditional NLP classifiers are not effective against LLM-based models (Carlini et al., 2023). Therefore, we use PromptAttack (Xu et al., 2023), a recent approach designed specifically to fool LLM-based classifiers. We consider a gray-box setting, where the adversary is unaware of the exact LLM used by the defender but has knowledge of the prompts employed for toxicity classification. The adversary has no query access to the victim classifier.
@@
Given a context-response sample $x$, PromptAttack uses an off-the-shelf LLM, $LLM_{p}$, to perturb the sample into an adversarial sample $x_{adv}$. This is done using a prompt-based strategy with three parts: (a) the input sample $x$, (b) an attack objective to fool a surrogate LLM-based toxicity classifier $LLM_{s}$ while maintaining semantic similarity between $x$ and $x_{adv}$, and (c) attack guidance with 9 different strategies at the character, word, and sentence levels. PromptAttack recommends using the same model for $LLM_{p}$ and $LLM_{s}$. It generates multiple candidates via an ensemble strategy and selects the one that successfully fools $LLM_{s}$ with the highest semantic similarity. We use LLaMA-2-Chat 13B as $LLM_{p}$ and 7B as $LLM_{s}$. [^4] LLaMA-2-Chat 7B gave better attack performance than 13B model. The prompt template is in Figure 5 in Appendix.
@@
Optimus (NH)  0.2 6.17 0.109   1.8 6.23 0.103
@@
Table 6 shows results for Optimus using the refusal classifier (NH setting) on the challenging LLaMA-2 chatbot. For the offensive and specialized categories, our adversarial samples successfully inject toxicity (RTR $47\%$), while preserving semantic similarity. Our Optimus defense effectively mitigates toxicity in both categories (RTR $1.8\%$), significantly lower than no-attack. Upon further investigation, we find that PromptAttack did not produce highly effective adversarial samples. For example, in offensive, there was only a 2.86% degradation in Recall ($\Delta R$). We observe similar outcomes for CH setting (results omitted for brevity).
7.1.2.. Jailbreak attacks
Jailbreak attacks bypass LLM safety mechanisms to generate otherwise restricted responses (Zhang et al., 2025). We leverage these attacks to fool our refusal classifier and investigate two types: (1) Manually-designed and (2) Optimization-based. We consider a white-box attack where the adversary has full knowledge and access to the classifier. However, the attacker can only add jailbreak prompts to context-response pairs (i.e., poisoned samples) and cannot modify the prompt used by our refusal classifier.
[…]
Optimization-based jailbreak attack.    This attack optimizes an adversarial suffix $suffix_{adv}$ that when appended to a context-response pair $x$ is classified as non-toxic by the refusal classifier. We use Liu et al. (Liu et al., 2024b)’s approach to craft a universal adversarial suffix tailored to a specific instruction (i.e., misclassify toxic samples). This approach is an improvement over the Greedy Coordinate Gradient (GCG) jailbreak attack. Optimization requires a surrogate LLM, which is the same as the victim classifier (white-box attack). Optimization is performed on 5 samples for 1K epochs with a batch size of 256. We use a token length of 150 for the adversarial suffix. Figure 6 (Appendix) shows an example attack.
While the optimization-based attack degrades our refusal classifier’s Recall by a significant 52.23%, Optimus still effectively mitigates toxicity, achieving a 7.9% RTR (well below the no-attack level). This is attributed to Optimus’s resilience to the degradation of the toxicity classifier’s performance (see analysis in Section 6.6).
@@
Optimus (CH)  4.5 5.85 0.099   7.9 5.78 0.101
[…]
7.2. Attack against healing data generation
@@
In this attack, the adversary tricks the safety-aligned LLM used by Optimus into producing toxic responses instead of contextually relevant healing responses. We aim to produce a different toxic response for each input context. However, the optimization-based attack in the previous section 7.1 is prohibitively expensive in this setting, To address this, we use AmpleGCG-Plus (Kumar et al., 2024)—an LLM trained on a large pool of adversarial suffixes—to efficiently generate tailored suffixes that can jailbreak target LLMs.
@@
We use AmpleGCG-Plus with a surrogate classifier that is identical to the victim classifier. For each context, we append a jailbreak instruction (e.g., ‘produce a harmful response’) and generate 10 suffix candidates using AmpleGCG-Plus. Figure 6 (Appendix) shows an example attack input. Suffixes are generated with a token length constrained between 20 and 50 tokens. We select the suffix yielding highest toxicity score for the response based on the surrogate classifier. When evaluated using the specialized category, 10.54% of adversarial contexts produce toxic responses, indicating a successful attack. These adversarial samples successfully inject toxicity leading to $50.06\%$ RTR. When Optimus is applied, the RTR reduces significantly to 17.3%, coming close to no-attack setting (13.1%). Note, this is a highly challenging, white-box attack that violates the integrity of the healing data used by our DPO scheme. Despite these challenges, Optimus is able to significantly mitigate toxicity.
7.3. Adversarial attacks in Dialog-based learning
We evaluate Optimus against adaptive adversarial attacks in a Dialog-based Learning (DBL) setting, as introduced by Weeks et al. (Weeks et al., 2023). In this post-deployment scenario, a chatbot is periodically retrained on recent user interactions. Attackers employ LLM-based agents to masquerade as normal users, engaging in toxic conversations to poison subsequent training cycles. The study defines two strategies: (1) Indiscriminate attacks, which generate toxic responses regardless of input; and (2) Backdoor attacks, which trigger toxicity only via specific phrases. We focus on the adaptive attacker aiming to bypass filters (i.e., toxicity classifiers), which serves as a robust evaluation for Optimus due to the attack’s practical relevance and distinct toxic language distribution.
We reproduce these attacks using the original code and datasets on a DD-BART base model (BART fine-tuned on DailyDialog). We implement Optimus using the refusal classifier and compare it against the Multi-level toxicity classifier from (Weeks et al., 2023). Table 8 details the RTR averaged over 5 trials. Optimus successfully mitigates both indiscriminate and backdoor attacks, reducing RTR to nearly 0 and outperforming the baseline Multi-level filter. Further experimental details are provided in Appendix A.3.